The crop package

Melchior FRANZ

May 20, 2003

Abstract

This article describes the crop package^{1}, which provides different forms of crop marks for trimming paper stacks, for camera alignment and for visualizing the page dimensions. There are options for centering the document page on the paper sheet, for mounting pages on a physical sheet, for reflecting and inverting the whole document or printing it upside-down, and for suppressing either text or graphics output.

The package was originally developed for needs of the Austrian Red Cross/Federal Province of Vienna/Department of Radiation Protection.

Contents

1 Introduction 1

2 How to use the package 2 2.1 Conventional options … 2 2.2 Runtime options . . . . . 3 2.3 Loading . . . . . . . . . . 5 2.4 Color support . . . . . . . 6 2.5 Custom document page

size . . . . . . . . . . . . . 6 2.6 Custom printing paper

sheet size . . . . . . . . . 6 2.7 Defining your own marks 7 2.8 The configuration file … 7

3 How the package works 8 3.1 The kernel mechanism . . 8

3.2 Dependencies . . . . . . . 8

4 The implementation 9 4.1 Preamble . . . . . . . . . 9 4.2 The device drivers …. 9 4.3 Size options . . . . . . . . 12 4.4 Runtime options handling 12 4.5 Axes and page info …. 13 4.6 The marks . . . . . . . . . 15 4.7 The kernel . . . . . . . . . 16 4.8 Mounting . . . . . . . . . 18 4.9 Page manipulation …. 19 4.10 Color handling . . . . . . 20 4.11 The graphics commands . 22 4.12 Final settings . . . . . . . 22 4.13 A mount4 example …. 23

1 Introduction

It is convenient to print documents for smaller logical paper sizes on paper of the printer’s standard physical paper size. On the one hand this keeps from changing paper stacks, on the other hand it allows printing close to the logical paper edge and even outside the logical page.

1This file has version number 1.9, last revised 2003/05/20.

I’d like to thank Rolf Niepraschk for his useful hints and suggestions, which influenced the package substantially. A big thank you also goes to Walter Schmidt for his extensive tests and his expertise on compatibility issues with different devices.

For trimming a whole paper stack or lining up the single pages on printing plates for photographical duplication a set of corner marks is required.

2 How to use the package

2.1 Conventional options

These options may only be used in the preamble and have to be stated as argu-

ments to the \usepackage command (as in \usepackage[mirror]{crop}).

a0, a1, a2, a3, a4, a5, a6, b0, b1, b2, b3, b4, b5, b6, letter, legal, executive These options declare the printing paper dimensions. One of them should be specified if the center option or one of the options dvips, pdftex and vtex is used. The size options do not define the logical document page size! See section 2.5 for how to achieve this.

width, height — Instead of using one of the pre-defined paper formats as described above, you can also set the physical paper dimensions directly. You may omit the ‘true’ specifier if you don’t plan to scale the document. Ex-

ample: \usepackage[cam,width=10truecm,height=13truecm]{crop}

center This option centers the logical document page on the physical printer paper and therefore requires that you declare the sheet size properly. Write,

for example, \usepackage[cam,a4,center]{crop} to center a document

of any size on ISO-A4 sheets.

landscape — Use this option in addition to the center option if you want to center a document on landscape oriented paper. It has nothing to do with L^{A}TEX’s landscape document class option.

dvips, pdftex, pdflatex, vtex, nodriver — If you are working with dvips, pdftex or vtex you may want to pass the dimensions of the paper that you are planning to print on to the respective driver program. Especially viewer programs like gs or gv make use of this bounding box information. Unfortunately, this can’t be done in a generic way—there’s no standard. These options select driver specific methods to set the paper size and to rotate and reflect a page. The crop package tries to find out by itself which driver to use. You find its choice mentioned in the log file. Additionally, you can suggest ([dvips]) or enforce ([dvips!]) one of the drivers. In the latter case you only have to add an exclamation point to the driver option. The difference is, that a suggestion may get overruled by the package. Assume you have asked for [dvips], but run the document file through pdflatex. In this case crop will automatically use the pdflatex driver. You can also force crop not to use any of the drivers by requesting the nodriver option. pdflatex is a synonym for pdftex.

mirror This option reflects the whole document, provided that the selected output driver supports the graphics package’s \reflectbox command. It doesn’t have any effect on the DVI file. rotate Rotates the document by 180^{∘} so that it appears upside-down. This may

be useful to circumvent problems with printers, which do not print close enough to the lower paper edge due to their paper feed mechanism.

invert Lets the whole document be printed white onto black background, if the color package can be loaded and the document is output with a color aware device. All further color changing commands stated in the document are disabled. This option doesn’t invert pictures, nor does it really swap text and paper color. Red text on green will still become white text on black. invert is stronger than notext.

notext — This option uses the color package to turn text to white color, after which all further color switching commands are disabled. This makes the text disappear from the printout, although it remains in the output file. See the description of the options nographics and graphics on page 4 for an explanation. This option is ignored if option invert was also requested.

2.2 Runtime options

These options may be used in the preamble like the ‘conventional’ options (see section 2.1), but also as optional arguments to the \crop command everywhere in the document (as in \crop[frame]). Using this command without options implies

\crop[cam,noaxes].

cam This mode provides four different marks (see figure 1), one for each corner. They indicate the logical paper edges without touching them and can thus be printed on every page. These marks are mainly thought for camera alignment. The \crop command selects this mode if no other mode is requested.

cross This mode provides four 4 cm wide crosses (see figure 1), one at each corner, that touch the logical paper edge. That’s why they should be printed on an extra page that will be used as a cover page while trimming the whole paper stack. (This is also the Red Cross mode ;-)

frame This mode draws a frame around the logical page and is mainly thought for visualizing the document page dimensions.

off This ‘option’ makes only sense in connection with the \crop command (i. e. at runtime). It disables all markings and is selected by default if the package is input without requesting any of the marks.

odd, even — Use these options to let the crop marks be put on odd/even pages only. They automatically turn on cam marks if no other marks have been requested. Note that only the page number is considered. If you have two subsequent pages both with page number 1, and ask for the odd option, then both pages will have marks.

axes, noaxes — These options enable/disable the output of little marks that indicate the logical page’s horizontal and vertical middle axis and may be selected in addition to one of the main modes. These marks might be needed for punching. Note that they are lost after trimming, since they lie outside the logical page. These marks are disabled by default. info, noinfo — Print the page information consisting of filename, date, time, page

number and page index on every sheet (see figure 1). The page index starts with #1 and is incremented with every page info line, hence being more reliable than page numbers, which are not unique and may be negative or contain letters. It can also be seen as a crop marks counter. Pages without crop marks aren’t counted. This page information is enabled by default.

font The page info line uses \normalfont by default. If you are typesetting the document in non-latin glyphs or a decorative, but less legible font, you may want to request a specific font for that info. Just assign a font switching command like \textsf to the font option parameter, leaving the initial backslash away: [font=textsf]. This command may take one argument (like \textsf{}) or stand alone (like \small). You can, of course, define a more complex command first, and assign that one:

\newcommand*\infofont[1]{\textcolor{blue}{\textsf{\small#1}}}
\crop[font=infofont]

color You can set the color of crop marks, axes and info text with this option, if the color package could be loaded. The option takes only color names, as in [color=red]. See the color package documentation for how to define custom colors.

mount1, mount2 — If more than one logical page is to be mounted on a physical sheet, you normally don’t want marks to appear on the inner edges, where the pages touch each other. The mount2 option prints only the outer marks. There’s also a mount1 option that is selected by default. These commands take a number as an optional argument serving as page offset. Type mount2 or mount2=0 for odd pages right and mount2=1 for odd pages left. Since further modes are likely to be document, driver, and printer dependent, it is up to you to implement them yourself. (See a mount4 suggestion on page 23.)

horigin, vorigin — The top and left margin are by default 1 inch wide. This can be changed using the dimensions \oddsidemargin, \evensidemargin and \topmargin. It’s more convenient, though, to let the geometry package define all these and further parameters. The options horigin and vorigin only move the marks and don’t change the page contents. Using these options is almost always a mistake, so use them only as a last resort! Both options take a (mandatory) dimension. These dimensions describe the way from the reference point—the upper left corner of the text block— to the upper left corner of the page in a Cartesian coordinate system. As both horigin and vorigin are by default −1 inch, you would for example write horigin=-.6in to move the marks by 0.4 inch to the right.

graphics, nographics — Color printouts are often more expensive than blackand-white ones, while their text quality is sometimes reduced. Therefore it may be desirable to create two versions of a document, one with only text and one with only graphics. Now you can feed the concerned pages to a color printer to print the notext version, and then to a mono laser printer with the nographics version. The graphics option turns graphics

[figure: “crop” — 2004/1/25 — 21:59 — page 5 — #1]

Figure 1: That’s what you see on top of a 9 cm wide document page when cam mode is requested: the marks, jobname, date, time, page number and crop marks index.

on again. You may want to mark up all colored pictures so that you can decide in the preamble, whether they shall be printed or not.

        %\newcommand*\colorgraphic{}                 % print them;
        \newcommand*\colorgraphic{\crop[nographics]} % don't!
        ...
        {\colorgraphic
        \includegraphics{...}
        \caption{...}
        }

2.3 Loading

Since all marks lie outside the logical page, the horizontal and vertical offset should be set properly. Otherwise the marks are likely to be cut off by the DVI driver or the printer. Provided that you have declared the size of your printing paper, you can use the center option to center every logical page on the respective sheet. There’s, however, no harm in centering an A4 page on A4 paper, in which case both offsets are set to 0 pt (unless, of course, you have set \mag ≠ 1000).

        \documentclass[a5paper]{article}
        \usepackage[cam,a4,center]{crop}
        \begin{document}
        ...
        \end{document}

                                           \documentclass[a5paper]{article}
                                           \usepackage[a4,center]{crop}
                                           \begin{document}
                                           ...
                                           \crop  % or: \crop[cross], etc.
                                           ...
                                           \end{document}

You get corner markings at every page shipped out after a cam, cross, or frame \crop mode request until you turn them off by typing \crop[off], or the actual

grouping level ends. Typing \crop without argument is equivalent to typing \crop[cam,noaxes]. Axis marks appear only together with one of the modes as listed above. If you only want one cover page for trimming, make sure that a page is actually output in the scope of \crop, for example:

    \newpage
    {\crop[cross,axes]\mbox{}\newpage}

2.4 Color support

The crop package always tries to load the color package. You can change the color of the physical page as usual by using the \pagecolor command in the preamble. But after that, within the document environment, \pagecolor is redefined to only color the logical page. The color of crop marks, axes marks and page info can independently be set via the color option. The options invert and notext override any color settings.

2.5 Custom document page size

The crop package respects any page layout that you specify by means of L^{A}TEX dimensions. The following example uses the geometry package, which I strongly recommend. Let’s assume you want to print a CD booklet (4^{23}/_{32} × 4^{3}/_{4} inch) on ISO-A4 paper:

    \documentclass{article}
    \usepackage[dvips=false,pdftex=false,vtex=false]{geometry}
    \geometry{
       paperwidth=4.71875in,
       paperheight=4.75in,
       margin=2em,
       bottom=1.5em,
       nohead
    }
    \usepackage[cam,a4,center,dvips]{crop}
    \begin{document}
    ...
    \end{document}

Note that the crop package should always be requested after setting up the ‘geometry’. See the geometry documentation for details. Always disable all of geometry’s driver options. While this isn’t necessary in every case, it doesn’t hurt and it makes your document more portable. You never know how the local geometry.cfg file on other workstations looks like!

2.6 Custom printing paper sheet size

If you want to use one of the center, dvips, pdftex or vtex options together with non-standard printing paper, you can set it via the width and height option, or simply add the respective paper definition to your crop.cfg file (see 2.8). Let’s for example define a new weird paper format, whereby the first dimension shall describe the paper width. Don’t forget to request true dimensions, otherwise you will get really weird results with scaled documents.

    \DeclareOption{weird}{\CROP@size{12truecm}{34truecm}}

Now you can use your new printing paper format like the pre-defined ones.

    \usepackage[frame,weird,center]{crop}

If you don’t need that format regularly or don’t want to depend on a crop.cfg file, then you might prefer to declare the dimensions in the document:

    \usepackage[frame,width=12truecm,height=34truecm,center]{crop}

2.7 Defining your own marks

If you need a funny mode, you can easily define it with only a couple of macros.

\cropdef The \cropdef command defines the mode switch. It takes as arguments: the name of a macro providing the page info (optional; enclosed in brackets), four macro names to be assigned to the upper left, the upper right, the lower left, and the lower right corner, each representing a picture with zero width and height, or \relax, and finally the mode name. The optional brackets may also be empty, if no page info is wanted, or contain the info code instead of a macro name.

    \newcommand*\funnymarkA{%         % a little x
      \begin{picture}(0,0)
        \thinlines\unitlength1pt
        \put(-5,-5){\line(1,1){10}}
        \put(-5,5){\line(1,-1){10}}
      \end{picture}}
    \newcommand*\funnymarkB{%         % a bullet
      \begin{picture}(0,0)
        \unitlength1pt
        \put(0,0){\circle*{5}}
      \end{picture}}
    \newcommand*\funnyinfo{funny page info}
    \cropdef[\funnyinfo]\relax\funnymarkA\relax\funnymarkB{funny}

Now you can select your new mode by typing \crop[funny].

Each of the axis marks is a picture that you can easily replace by some custom definition. There’s no setup command like \cropdef, though. The kernel provides two ‘hooks’ that can be used to add local extensions. These are macros that default to \relax. The first, \CROP@user@a, is executed at every page, no matter if marks are shown or not, while the second, \CROP@user@b is only executed at pages that contain crop marks. Local definitions and modifications are ideally put into a local configuration file:

2.8 The configuration file

If you want to change the predefined settings or add new features, then create a file named ‘crop.cfg’ and put it in a directory, where TEX can find it. This configuration file will then be loaded at the end of the crop.sty file, so you can redefine any settings or commands therein, select package options and even introduce new ones. But if you intend to give your documents to others, don’t forget to give them the required configuration files, too! That’s how such a file could look like:

    % define a new printing paper size
    \DeclareOption{special}{\CROP@size{22truecm}{37truecm}}

    % make the internal time string (used in the page
    % information) accessible in the whole document
    \let\Time\CROP@time
    % let's use a different font for the predefined page
    % information (we could also have written
    % \newcommand*\CROP@font[1]{\textsf{#1}})
    \crop[font=textsf]
    \endinput

3 How the package works

3.1 The kernel mechanism

TEX outputs a page via the \shipout command. The crop package redefines \shipout to insert the requested marks before it outputs the page contents. It is carefully designed to coexist peacefully with other packages, which use the same method (like the everyshi package by Martin Schröder, from whom I have in fact borrowed some ideas).

In addition to the crop marks every page gets an info line containing the jobname, the current date and time, the page number and an index number printed on top. This line can be turned off (noinfo) and on (info) anywhere in the document.

3.2 Dependencies

3.2.1 latex.ltx

The package works with all L^{A}TEX 2ε standard classes (tested with L^{A}TEX 2ε 1997/12/01 and sporadically with later versions), it does not work with plain TEX.

The crop package uses (and relies on) the internal L^{A}TEX tokens \hb@xt@,

\filename@parse, \@classoptionslist, \@ifundefined, \@height, \@depth,
\filename@base \@width, \z@, \@ne, \z@skip, \p@, \c@page, \@namedef,
\@nameuse, \strip@pt, \two@digits, \count@, \dimen@, \@for, \@empty,

\@gobble and \@undefined, all of which are expected to keep their current meaning in future L^{A}TEX 2ε releases. The crop package will, however, be supported at least for some years, so you needn’t worry about it.

3.2.2 color.sty

crop’s color handling depends on the color package. The following inter-

nal macros are used directly: \@declaredcolor, \current@color, \set@color,

\set@page@color (Tested with color.sty, version 1.0i as of 1999/02/16.)

3.2.3 graphics.sty

crop’s driver detection, as well as the options rotate mirror, and nographics depend on the graphics package. The following internal macros are used directly:

\Gin@PS@raw, \Ginclude@graphics, \Gin@driver (Tested with graphics.sty,

version 1.0l as of 1999/02/16.)

4 The implementation

4.1 Preamble

\stockwidth \stockheight \CROP@index \CROP@font Make sure that \stockwidth and \stockheight are \dimen registers that hold

the physical paper size. They are initially set to the paper size, but will be changed by the size options. These registers are also used/provided by the memoir class and the hyperref package. The \CROP@font macro is by default empty and can be changed through the font option.

1 ⟨∗package⟩

 2 \NeedsTeXFormat{LaTeX2e}
 3 \ProvidesPackage{crop}[2003/05/20 v1.9 crop marks  (mf)]
 4 \expandafter\ifx\csname stockwidth\endcsname\relax
 5  \newdimen\stockwidth
 6  \stockwidth\paperwidth
 7 \fi
 8 \expandafter\ifx\csname stockheight\endcsname\relax
 9  \newdimen\stockheight
10  \stockheight\paperheight
11 \fi
12 \newcount\CROP@index
13 \CROP@index\z@
14 \newcommand*\CROP@font{}

\CROP@stockcolor \CROP@pagecolor \CROP@needscolor Try to load the color package. It is needed for the invert and the notext option and, of course, for the modified \pagecolor command. Changing the meaning of \current@color looks dangerous, but it is only done if the color package couldn’t be loaded, anyway.

15 \let\CROP@stockcolor\@empty
16 \let\CROP@pagecolor\@empty
17 \IfFileExists{color.sty}{%
18  \RequirePackage{color}%
19  \let\CROP@needscolor\@empty
20 }{%
21  \newcommand*\CROP@needscolor{%
22  \PackageError{crop}{%
23  The `invert' and `notext' options require the\MessageBreak
24  `color' package, which doesn't seem to be installed%
25  }{%
26  Install the `color' package or don't use the `invert'
27  \MessageBreak or `notext' option.
28  }%
29  \let\CROP@needscolor\relax
30  }%
31  \let\current@color\relax

32 }

4.2 The device drivers

\CROP@detdriver \CROP@Ginclude@graphics \CROP@ps The options graphics and nographics depend on the graphics package, which, if configured appropriately, also tells us which output device is preferred on the system. Show a warning, if the package couldn’t be loaded, because we have to use a less portable PS ‘driver’ then.

33 \let\CROP@detdriver\@empty
34 \IfFileExists{graphics.sty}{%
35  \RequirePackage{graphics}%
36  \let\CROP@Ginclude@graphics\Ginclude@graphics
37  \ifx\Gin@driver\@empty\else
38  \filename@parse{\Gin@driver}%
39  \edef\CROP@detdriver{\filename@base}%
40  \fi
41  \let\CROP@ps\Gin@PS@raw
42 }{%
43  \PackageWarning{crop}{I couldn't find the `graphics' package, so
44  I'll use\MessageBreak my internal PostScript interface%
45  }%
46  \newcommand*\CROP@ps[1]{\special{ps: ##1}}%

47 }

\CROP@reqdriver Define options that suggest …

48 \let\CROP@reqdriver\@empty
49 \DeclareOption{vtex}{\def\CROP@reqdriver{vtex}}
50 \DeclareOption{pdftex}{\def\CROP@reqdriver{pdftex}}
51 \DeclareOption{pdflatex}{\def\CROP@reqdriver{pdftex}}
52 \DeclareOption{dvips}{\def\CROP@reqdriver{dvips}}

\CROP@driver … or enforce a graphics driver. Note the exclamation points!

53 \let\CROP@driver\@empty
54 \DeclareOption{vtex!}{\def\CROP@driver{vtex}}
55 \DeclareOption{pdftex!}{\def\CROP@driver{pdftex}}
56 \DeclareOption{pdflatex!}{\def\CROP@driver{pdftex}}
57 \DeclareOption{dvips!}{\def\CROP@driver{dvips}}
58 \DeclareOption{nodriver}{\def\CROP@driver{none}}
59 \DeclareOption{!}{\def\CROP@driver{none}}
\CROP@evaldriver If \CROP@driver wasn’t already set, decide \AtBeginDocument which graphics

driver to use. A detected driver takes precedence over a ‘suggested’ one. Show a warning if the user’s choice is ignored.

60 \newcommand*\CROP@evaldriver{%
61  \ifx\CROP@driver\@empty
62  \PackageInfo{crop}{requested driver: `\CROP@reqdriver'}%
63  \ifx\pdfoutput\@undefined\else
64  \ifx\pdfoutput\relax\else
65  \ifcase\pdfoutput\else
66  \def\CROP@detdriver{pdftex}%
67  \fi
68  \fi
69  \fi
70  \ifx\VTeXversion\@undefined\else
71  \ifx\VTeXversion\relax\else
72  \def\CROP@detdriver{vtex}%
73  \fi
74  \fi
75  \PackageInfo{crop}{detected driver: `\CROP@detdriver'}%

76  \ifx\CROP@reqdriver\@empty\else
77  \ifx\CROP@reqdriver\@empty\else
78  \ifx\CROP@reqdriver\CROP@detdriver\else
79  \PackageWarningNoLine{crop}{%
80  You requested the `\CROP@reqdriver' driver
81  but I think that\MessageBreak the
82  `\CROP@detdriver' driver works better in the
83  current\MessageBreak context. You can force
84  me to respect your decision\MessageBreak
85  by adding an exclamation point as in
86  [\CROP@reqdriver!]%
87  }%
88  \fi
89  \fi
90  \fi
91  \ifx\CROP@detdriver\@empty
92  \let\CROP@driver\CROP@reqdriver
93  \else
94  \let\CROP@driver\CROP@detdriver
95  \fi
96  \fi
97  \let\CROP@evaldriver\relax

98 }

99 \AtBeginDocument{\CROP@evaldriver}
\CROP@init@dvips \CROP@init@pdftex \CROP@init@vtex \CROP@init@none These macros prepare the crop package for one of the supported graphics drivers.

They don’t do anything spectacular, they just select the proper rotate and mirror macro and hand the physical paper size over to the driver program.

100 \newcommand*\CROP@init@dvips{%
101  \PackageInfo{crop}{using `dvips' graphics driver}%
102  \AtBeginDvi{%
103  \special{papersize=\the\stockwidth,\the\stockheight}%

104 }% 105 }

106 \newcommand*\CROP@init@pdftex{%
107  \PackageInfo{crop}{using `pdftex' graphics driver}%
108  \pdfpagewidth\stockwidth
109  \pdfpageheight\stockheight
110  \let\CROP@reflect\CROP@genreflect
111  \let\CROP@rotate\CROP@genrotate

112 }

113 \newcommand*\CROP@init@vtex{%
114  \PackageInfo{crop}{using `vtex' graphics driver}%
115  \mediawidth\stockwidth
116  \mediaheight\stockheight
117  \let\CROP@reflect\CROP@genreflect
118  \let\CROP@rotate\CROP@genrotate

119 }

120 \newcommand*\CROP@init@none{%
121  \PackageInfo{crop}{not using any graphics driver}%

122 }

4.3 Size options

\CROP@size \CROP@opt@width \CROP@opt@height These options set different standard printing paper sizes, which are needed for centering and as a hint for the dvips, pdftex or vtex program. Since the physical paper dimensions must not underlie a possible scaling, true dimensions are used. The landscape option exchanges the \hoffset and \voffset values.

123 \newcommand*\CROP@size[2]{\stockwidth#1 \stockheight#2 }
124 \DeclareOption{landscape}{%
125  \def\CROP@size#1#2{\stockheight#1 \stockwidth#2 }%

126 }

127 \DeclareOption{a0}{\CROP@size{841truemm}{1189truemm}}
128 \DeclareOption{a1}{\CROP@size{595truemm}{841truemm}}
129 \DeclareOption{a2}{\CROP@size{420truemm}{595truemm}}
130 \DeclareOption{a3}{\CROP@size{297truemm}{420truemm}}
131 \DeclareOption{a4}{\CROP@size{210truemm}{297truemm}}
132 \DeclareOption{a5}{\CROP@size{149truemm}{210truemm}}
133 \DeclareOption{a6}{\CROP@size{105truemm}{149truemm}}
134 \DeclareOption{b0}{\CROP@size{1000truemm}{1414truemm}}
135 \DeclareOption{b1}{\CROP@size{707truemm}{1000truemm}}
136 \DeclareOption{b2}{\CROP@size{500truemm}{707truemm}}
137 \DeclareOption{b3}{\CROP@size{353truemm}{500truemm}}
138 \DeclareOption{b4}{\CROP@size{250truemm}{353truemm}}
139 \DeclareOption{b5}{\CROP@size{176truemm}{250truemm}}
140 \DeclareOption{b6}{\CROP@size{125truemm}{176truemm}}
141 \DeclareOption{letter}{\CROP@size{8.5truein}{11truein}}
142 \DeclareOption{legal}{\CROP@size{8.5truein}{14truein}}
143 \DeclareOption{executive}{\CROP@size{7.25truein}{10.5truein}}
144 \newcommand\CROP@opt@width{\stockwidth\CROP@@}
145 \newcommand\CROP@opt@height{\stockheight\CROP@@}

\CROP@center The center option sets \voffset and \hoffset so that the document pages are centered on the printing paper sheets.

146 \DeclareOption{center}{\AtBeginDocument{\CROP@center}}
147 \newcommand*\CROP@center{%
148  \voffset\stockheight
149  \advance\voffset-\paperheight
150  \voffset.5\voffset
151  \hoffset\stockwidth
152  \advance\hoffset-\paperwidth
153  \hoffset.5\hoffset

154 }

4.4 Runtime options handling

Pass every unknown option to the macro \CROP@execopt.

155 \DeclareOption*{\CROP@execopt\CurrentOption}

\crop The \crop macro allows options to be used both in the preamble and throughout the document. Every argument of the optional argument list is passed to the macro \CROP@execopt. The options cam and noaxes are selected by default.

156 \newcommand*\crop[1][cam,noaxes]{%
157  \@for\CROP@@:=#1\do{\CROP@execopt\CROP@@}%

158 }

\CROP@execopt Every execution of this macro with an argument n leads to the execution of a macro \CROP@opt@n or a warning if no such exists. Optional arguments (separated by an equal sign) are cut off and stored in \CROP@@. The macro tolerates even arguments for options that are not prepared to handle arguments (e. g. cross=garbage), or more than one argument (e. g. mount2=1=garbage). This makes the design simpler and doesn’t hurt.

159 \newcommand*\CROP@execopt[1]{%
160  \def\CROP@##1=##2=##3\@nil{\def\CROP@{##1}\def\CROP@@{##2}}%
161  \expandafter\CROP@#1==\@nil%
162  \@ifundefined{CROP@opt@\CROP@}{%
163  \PackageError{crop}{%
164  Requested option `#1' not provided%
165  }{%
166  Note that the `*center' options are obsolete. You have to
167  request\MessageBreak e.g. [a4,center] instead of
168  [a4center].

169 }%

170  }{%
171  \@nameuse{CROP@opt@\CROP@}%

172 }% 173 }

\cropdef The \cropdef macro defines a mode switch (see section 2.7). It supports only corner marks and the page info, but not the axis marks, mainly for hysterical raisins.

174 \newcommand*\cropdef[6][\CROP@@info]{%
175  \@namedef{CROP@opt@#6}{%
176  \def\CROP@info{#1}%
177  \let\CROP@ulc#2
178  \let\CROP@urc#3
179  \let\CROP@llc#4
180  \let\CROP@lrc#5
181  \let\CROP@@@marks\CROP@marks

182 }% 183 }

4.5 Axes and page info

\CROP@@laxis \CROP@@raxis \CROP@@upaxis \CROP@@loaxis The standard definitions for the axes option.
184 \newcommand*\CROP@@laxis{%
185  \begin{picture}(0,0)
186  \unitlength\p@\thinlines
187  \put(-2,0){\line(-1,0){11}}
188  \end{picture}%

189 }

190 \newcommand*\CROP@@raxis{%
191  \begin{picture}(0,0)
192  \unitlength\p@\thinlines
193  \put(2,0){\line(1,0){11}}
194  \end{picture}%

195 }

196 \newcommand*\CROP@@upaxis{%

197  \begin{picture}(0,0)
198  \unitlength\p@\thinlines
199  \put(0,2){\line(0,1){11}}
200  \end{picture}%

201 }

202 \newcommand*\CROP@@loaxis{%
203  \begin{picture}(0,0)
204  \unitlength\p@\thinlines
205  \put(0,-2){\line(0,-1){11}}
206  \end{picture}%

207 }

\CROP@time \CROP@@info \CROP@opt@font This macro prints the jobname, the current date and time, the page number and an index number at the top of the logical page.

208 \newcommand*\CROP@time{}
209 \bgroup
210  \count@\time
211  \divide\time60
212  \count\@ne\time
213  \multiply\time60
214  \advance\count@-\time
215  \xdef\CROP@time{\the\count\@ne:\two@digits{\count@}}
216 \egroup
217 \newcommand*\CROP@@info{{%
218  \global\advance\CROP@index\@ne
219  \def\x{\discretionary{}{}{\hbox{\kern.5em---\kern.5em}}}%
220  \advance\paperwidth-20\p@
221  \dimen@4pt
222  \ifx\CROP@pagecolor\@empty
223  \else
224  \advance\dimen@\CROP@overlap
225  \fi
226  \hb@xt@\z@{%
227  \hss
228  \vbox to\z@{%
229  \centering
230  \hsize\paperwidth
231  \vss
232  \normalfont
233  \normalsize
234  \expandafter\csname\CROP@font\endcsname{%
235  ``\jobname''\x
236  \the\year/\the\month/\the\day\x
237  \CROP@time\x
238  page\kern.5em\thepage\x
239  \#\the\CROP@index
240  \strut

241 }%

242  \vskip\dimen@

243 }%

244  \hss

245 }% 246 }}

247 \newcommand*\CROP@opt@font{\let\CROP@font\CROP@@}

4.6 The marks

The following four macros provide different marks for the cam mode. They do not touch the logical page and can, thus, be printed on every single sheet.

\CROP@@ulc The cam mode corner mark for the upper left corner.

248 \newcommand*\CROP@@ulc{%
249  \begin{picture}(0,0)
250  \unitlength\p@\thinlines
251  \put(-30,0){\circle{10}}
252  \put(-30,-5){\line(0,1){10}}
253  \put(-35,0){\line(1,0){30}}
254  \put(0,30){\circle{10}}
255  \put(-5,30){\line(1,0){10}}
256  \put(0,35){\line(0,-1){30}}
257  \end{picture}%

258 }

\CROP@@urc The cam mode corner mark for the upper right corner.

259 \newcommand*\CROP@@urc{%
260  \begin{picture}(0,0)
261  \unitlength\p@\thinlines
262  \put(30,0){\circle{10}}
263  \put(30,-5){\line(0,1){10}}
264  \put(35,0){\line(-1,0){30}}
265  \put(0,30){\circle{10}}
266  \put(-5,30){\line(1,0){10}}
267  \put(0,35){\line(0,-1){30}}
268  \end{picture}%

269 }

\CROP@@llc The cam mode corner mark for the lower left corner.

270 \newcommand*\CROP@@llc{%
271  \begin{picture}(0,0)
272  \unitlength\p@\thinlines
273  \put(-30,0){\circle{10}}
274  \put(-30,-5){\line(0,1){10}}
275  \put(-35,0){\line(1,0){30}}
276  \put(0,-30){\circle{10}}
277  \put(-5,-30){\line(1,0){10}}
278  \put(0,-35){\line(0,1){30}}
279  \end{picture}%

280 }

\CROP@@lrc The cam mode corner mark for the lower right corner.

281 \newcommand*\CROP@@lrc{%
282  \begin{picture}(0,0)
283  \unitlength\p@\thinlines
284  \put(30,0){\circle{10}}
285  \put(30,-5){\line(0,1){10}}
286  \put(35,0){\line(-1,0){30}}
287  \put(0,-30){\circle{10}}
288  \put(-5,-30){\line(1,0){10}}

289  \put(0,-35){\line(0,1){30}}
290  \end{picture}%

291 }

\CROP@opt@cam Define the cam mode switch with four different marks.

292 \cropdef\CROP@@ulc\CROP@@urc\CROP@@llc\CROP@@lrc{cam}

\CROP@@cross This macro provides a 4 cm wide cross.

293 \newcommand*\CROP@@cross{%
294  \begin{picture}(0,0)
295  \unitlength1cm\thinlines
296  \put(-2,0){\line(1,0){4}}
297  \put(0,-2){\line(0,1){4}}
298  \end{picture}%

299 }

\CROP@opt@cross Define the cross mode switch with four times the same mark.

300 \cropdef\CROP@@cross\CROP@@cross\CROP@@cross\CROP@@cross{cross}

\CROP@@frame The frame mode draws a simple frame around the logical page. The frame mark is designed to be used in the upper left corner. Since graphics commands expect numbers without dimensions, \paperwidth and -height are transformed to numbers (representing printer’s points). This is done by stripping off the unit pt.

301 \newcommand*\CROP@@frame{%
302  \begin{picture}(0,0)
303  \unitlength\p@\thinlines
304  \put(0,0){\line(1,0){\strip@pt\paperwidth}}
305  \put(0,0){\line(0,-1){\strip@pt\paperheight}}
306  \put(\strip@pt\paperwidth,0){\line(0,-1){\strip@pt\paperheight}}
307  \put(0,-\strip@pt\paperheight){\line(1,0){\strip@pt\paperwidth}}
308  \end{picture}%

309 }

\CROP@opt@frame Define the frame mode switch with only one mark. The other corners may \relax.

310 \cropdef\CROP@@frame\relax\relax\relax{frame}

4.7 The kernel

\CROP@shipout \shipout \CROP@ship \CROP@shiplist \CROP@@ship These macros redefine the TEX primitive \shipout to insert the contents of the

macro \CROP@shiplist on top of the box which contains the page contents ready for output, after which the original \shipout command is executed.

311 \let\CROP@shipout\shipout
312 \renewcommand*\shipout{%
313  \afterassignment\CROP@ship
314  \setbox\@cclv=%

315 }

316 \newcommand*\CROP@ship{%
317  \ifvoid\@cclv
318  \expandafter\aftergroup
319  \fi
320  \CROP@@ship

322 \newcommand*\CROP@shiplist{%
323  \lineskip\z@
324  \lineskiplimit\z@
325  \baselineskip\z@
326  \CROP@kernel
327  \box\@cclv

328 }

329 \newcommand*\CROP@@ship{%
330  \CROP@shipout\vbox{\CROP@shiplist}%

331 }

\CROP@shipadd This macro adds a page manipulation command to the shiplist, which gets every ready page as argument.

332 \newcommand*\CROP@shipadd[1]{%
333  \bgroup
334  \toks@\expandafter{\expandafter#1\expandafter{\CROP@shiplist}}%
335  \xdef\CROP@shiplist{\the\toks@}%
336  \egroup

337 }

\CROP@kernel \CROP@marks \CROP@@@marks \CROP@setmarkcolor \CROP@user@a \CROP@user@b \CROP@opt@horigin \CROP@opt@vorigin  \CROP@kernel essentially contains a \vbox with zero width and height. The

\CROP@hook command—which normally equals \relax—allows to insert commands that modify the behavior of the selected mode (see the options mount1 and mount2). \CROP@user@a and \CROP@user@b are user definable hooks.

338 \newcommand*\CROP@kernel{%
339  \vbox to\z@{%
340  \vskip\CROP@vorigin
341  \hb@xt@\z@{%
342  \hskip\CROP@horigin
343  \vbox to\paperheight{%
344  \let\protect\relax
345  \hsize\paperwidth
346  \CROP@hook
347  \CROP@user@a
348  \CROP@drawstockcolor
349  \CROP@drawpagecolor
350  \CROP@@@marks

351 }%

352  \hss

353 }%

354  \vss

355 }% 356 }

357 \newcommand*\CROP@marks{%
358  \CROP@setmarkcolor
359  \CROP@user@b
360  \CROP@ulc\null\hfill\CROP@@@info\CROP@upedge\hfill\null\CROP@urc
361  \vfill
362  \CROP@ledge\hfill\CROP@redge
363  \vfill
364  \CROP@llc\null\hfill\CROP@loedge\hfill\null\CROP@lrc

365 }

366 \let\CROP@@@marks\CROP@marks
367 \newcommand*\CROP@setmarkcolor{%
368  \let\current@color\CROP@markcolor
369  \set@color

370 }

371 \let\CROP@user@a\relax
372 \let\CROP@user@b\relax
373 \newcommand*\CROP@opt@horigin{\let\CROP@horigin\CROP@@}
374 \newcommand*\CROP@opt@vorigin{\let\CROP@vorigin\CROP@@}

\CROP@opt@off \CROP@opt@odd \CROP@opt@even These macros start and stop the output of crop marks.

375 \newcommand*\CROP@opt@off{%
376  \let\CROP@@@marks\vfil

377 }

378 \newcommand*\CROP@opt@odd{%
379  \def\CROP@@@marks{\ifodd\c@page\CROP@marks\else\vfil\fi}%

380 }

381 \newcommand*\CROP@opt@even{%
382  \def\CROP@@@marks{\ifodd\c@page\vfil\else\CROP@marks\fi}%

383 }

\CROP@@@info \CROP@opt@info \CROP@opt@noinfo \CROP@opt@axes \CROP@opt@noaxes Enable and disable the output of axis marks and page info.
384 \newcommand*\CROP@@@info{}
385 \newcommand*\CROP@opt@info{\def\CROP@@@info{\CROP@info}}
386 \newcommand*\CROP@opt@noinfo{\let\CROP@@@info\relax}
387 \newcommand*\CROP@opt@axes{%
388  \let\CROP@ledge\CROP@@laxis
389  \let\CROP@redge\CROP@@raxis
390  \let\CROP@upedge\CROP@@upaxis
391  \let\CROP@loedge\CROP@@loaxis

392 }

393 \newcommand*\CROP@opt@noaxes{%
394  \let\CROP@ledge\relax
395  \let\CROP@redge\relax
396  \let\CROP@upedge\relax
397  \let\CROP@loedge\relax

398 }

4.8 Mounting

\CROP@opt@mount1 \CROP@opt@mount2 Since \newcommand doesn’t allow macro names to contain non-letters, we need a

 construction with \csname, \endcsname, and \expandafter. \@namedef would

have worked, too, but it would not have made a check for redefinitions.

399 \expandafter\newcommand\expandafter*\csname CROP@opt@mount1\endcsname{%
400  \let\CROP@hook\relax

401 }

402 \newcount\CROP@offset
403 \expandafter\newcommand\expandafter*\csname CROP@opt@mount2\endcsname{%
404  \CROP@offset=\ifx\CROP@@\@empty\z@\else\CROP@@\fi
405  \def\CROP@hook{%
406  \count@\c@page
407  \advance\count@\CROP@offset
408  \ifodd\count@

409  \let\CROP@ulc\relax
410  \let\CROP@llc\relax
411  \let\CROP@ledge\relax
412  \else
413  \let\CROP@urc\relax
414  \let\CROP@lrc\relax
415  \let\CROP@redge\relax
416  \fi

417 }% 418 }

4.9 Page manipulation

\CROP@reflect \CROP@genreflect \CROP@rotate \CROP@genrotate The mirror and rotate options add a macro to the shiplist, which then gets every output page and embeds it in a PostScript environment (dvips) or lets the graphics package reflect or rotate it (pdftex). We could also use the generic operations \CROP@genreflect and \CROP@genrotate for the dvips mode. They would produce correct PS documents, the intermediate DVI document, however, would be unreadable.

419 \DeclareOption{mirror}{%
420  \AtBeginDocument{\CROP@shipadd\CROP@reflect}

421 }

422 \newcommand*\CROP@reflect[1]{%
423  \vbox to\z@{%
424  \vskip\CROP@vorigin
425  \hb@xt@\z@{%
426  \hskip\CROP@horigin
427  \CROP@ps{gsave currentpoint}%
428  \kern\paperwidth
429  \CROP@ps{currentpoint}%
430  \hss

431 }%

432  \vss

433 }%

434  \CROP@ps{translate -1 1 scale neg exch neg exch translate}%
435  \vbox{#1}%
436  \CROP@ps{grestore}%

437 }

438 \newcommand*\CROP@genreflect[1]{%
439  \leavevmode
440  \dimen0\CROP@horigin
441  \kern2\dimen0
442  \reflectbox{%
443  \hb@xt@\paperwidth{%
444  \vbox to\paperheight{%
445  #1%
446  \vss

447 }%

448  \hss

449 }% 450 }% 451 }

452 \DeclareOption{rotate}{%

453  \AtBeginDocument{\CROP@shipadd\CROP@rotate}

454 }

455 \newcommand*\CROP@rotate[1]{%
456  \hb@xt@\z@{%
457  \hskip\CROP@horigin
458  \vbox to\z@{%
459  \vskip\CROP@vorigin
460  \CROP@ps{gsave currentpoint}%
461  \kern\paperheight
462  \hb@xt@\z@{%
463  \kern\paperwidth
464  \CROP@ps{currentpoint}%
465  \hss

466 }%

467  \vss

468 }%

469  \hss

470 }%

471  \CROP@ps{translate 180 rotate neg exch neg exch translate}%
472  \vbox{#1}%
473  \CROP@ps{grestore}%

474 }

475 \newcommand*\CROP@genrotate[1]{%
476  \dimen0\CROP@vorigin
477  \kern2\dimen0
478  \leavevmode
479  \dimen0\CROP@horigin
480  \kern2\dimen0
481  \rotatebox{180}{%
482  \hb@xt@\paperwidth{%
483  \vbox to\paperheight{%
484  #1%
485  \vss

486 }%

487  \hss

488 }% 489 }% 490 }

4.10 Color handling

\CROP@stockcolor \CROP@pagecolor \set@page@color \CROP@needscolor \CROP@defmarkcolor \CROP@opt@color \CROP@drawstockcolor \CROP@drawpagecolor \CROP@overlap \CROP@opt@overlap These macros care for the color of crop marks and of the logical and the physical

page. The overlap value is the amount that the logical page is drawn over the page boundaries on each side. This is necessary to get good results on imprecise cutting machines.

491 \newcommand*\CROP@defmarkcolor[1]{{%
492  \def\set@color{\global\let\CROP@markcolor\current@color}%
493  \@declaredcolor{#1}%

494 }}

495 \ifx\CROP@needscolor\@empty
496  \renewcommand*\set@page@color{%
497  \global\let\CROP@stockcolor\current@color

498 }%

499  \AtBeginDocument{%
500  \def\set@page@color{%
501  \global\let\CROP@pagecolor\current@color

502 }% 503 }%

504  \CROP@defmarkcolor{black}%
505  \let\CROP@needscolor\relax
506 \fi
507 \newcommand*\CROP@opt@color{%
508  \CROP@needscolor
509  \expandafter\CROP@defmarkcolor\expandafter{\CROP@@}%

510 }

511 \newcommand*\CROP@drawstockcolor{%
512  \ifx\CROP@stockcolor\@empty
513  \else
514  \rlap{%
515  \smash{%
516  \raise\voffset\hbox{%
517  \let\current@color\CROP@stockcolor
518  \set@color
519  \hskip-\hoffset
520  \vrule width\stockwidth height\z@ depth\stockheight

521 }% 522 }% 523 }%

524  \fi

525 }

526 \newcommand*\CROP@drawpagecolor{%
527  \ifx\CROP@pagecolor\@empty
528  \else
529  \rlap{%
530  \smash{%
531  \dimen@\CROP@overlap
532  \advance\paperwidth2\dimen@
533  \advance\paperheight2\dimen@
534  \raise\dimen@\hbox{%
535  \let\current@color\CROP@pagecolor
536  \set@color
537  \hskip-\dimen@
538  \vrule width\paperwidth height\z@ depth\paperheight

539 }% 540 }% 541 }%

542  \fi

543 }

544 \def\CROP@overlap{3truemm}
545 \newcommand*\CROP@opt@overlap{\let\CROP@overlap\CROP@@}

\CROP@invert The invert option simply switches to black background and white text, after which it disables all color switching commands. The notext option does the same with white text on white background. The \@gobble on the last line keeps notext from switching to white background and breaking a prior invert.

546 \newcommand*\CROP@invert[1]{%
547  \CROP@needscolor

548  \AtBeginDvi{%
549  \pagecolor{#1}%
550  \global\let\set@page@color\relax
551  \global\let\CROP@setpagecolor\relax

552 }%

553  \color{white}%
554  \DeclareRobustCommand*\color[2][]{}%
555  \let\pagecolor\color
556  \let\textcolor\color
557  \let\CROP@invert\@gobble

558 }

559 \DeclareOption{invert}{%
560  \CROP@invert{black}%
561  \let\CROP@setmarkcolor\relax

562 }

563 \DeclareOption{notext}{%
564  \CROP@invert{white}%

565 }

4.11 The graphics commands

\CROP@opt@nographics \CROP@opt@graphics The nographics option redefines the \Ginclude@graphics command from the

graphics package, so that it outputs its argument as a phantom. This makes the image invisible but takes up the same amount of white space. The graphics option re-enables graphics.

566 \newcommand*\CROP@opt@nographics{%
567  \def\Ginclude@graphics##1{%
568  \phantom{%
569  \CROP@Ginclude@graphics{##1}%

570 }% 571 }% 572 }%

573 \newcommand*\CROP@opt@graphics{%
574  \let\Ginclude@graphics\CROP@Ginclude@graphics

575 }

4.12 Final settings

\CROP@horigin \CROP@vorigin Switch off marks and axes, set one page per sheet, load the local configuration file,

and process the requested options. Finally: Exit.

Notice that we cannot simply use \ExecuteOptions to preselect options off, noaxes, info, and mount1, because it does not accept default options declared with \DeclareOption*. \@nameuse doesn’t complain if the command sequence is undefined. We let this only be executed \AtBeginDocument, because there are possibly commands from the center option in the queue that have to be processed first.

576 \newcommand*\CROP@horigin{-1truein}
577 \newcommand*\CROP@vorigin{-1truein}
578 \crop[cam,off,noaxes,info,mount1]
579 \InputIfFileExists{crop.cfg}{%
580  \PackageInfo{crop}{Local config file crop.cfg used}
581 }{}

2 1
0 3

Figure 2: Possible mount4 arrangement

582 \ProcessOptions
583 \AtBeginDocument{\@nameuse{CROP@init@\CROP@driver}}
584 \endinput

585 ⟨/package⟩

4.13 A mount4 example

Since a mount4 mode is likely to be subject to specific local needs, there’s only a suggestion provided, which supports a page arrangement as shown in figure 2.

First of all \CROP@offset is set to the value of the (optional) argument or zero. Then \CROP@hook is defined first to set \count@ to the page number increased by this offset: p = pagenumber + offset.

\expandafter\newcommand\expandafter*\csname CROP@opt@mount4\endcsname{%
\CROP@offset=\ifx\CROP@@\@empty\z@\else\CROP@@\fi
\def\CROP@hook{%
\count@\c@page
\advance\count@\CROP@offset

Now bits 0 and 1 are checked via \ifodd to get p modulo 4, after which the respective marks are deleted. The comments in the example use for simplicity C-notation in which ‘%’ is the modulo or remainder operator, ‘==’ the equal, and ‘||’ the logical (inclusive) OR operator.

\ifodd\count@                  %% if (p % 4 == 1 || p % 4 == 3)
\let\CROP@ulc\relax
\let\CROP@llc\relax
\let\CROP@ledge\relax
\divide\count@2
\ifodd\count@              %%    if (p % 4 == 3)
    \let\CROP@urc\relax
    \let\CROP@info\relax
    \let\CROP@upedge\relax
\else                      %%    if (p % 4 == 1)
    \let\CROP@lrc\relax
    \let\CROP@loedge\relax
\fi
\else                          %% if (p % 4 == 0 || p % 4 == 2)
\let\CROP@urc\relax
\let\CROP@lrc\relax
\let\CROP@redge\relax
\divide\count@2
\ifodd\count@              %%    if (p % 4 == 2)
    \let\CROP@llc\relax

\let\CROP@loedge\relax
\else                      %%    if (p % 4 == 0)
\let\CROP@ulc\relax
\let\CROP@info\relax
\let\CROP@upedge\relax
\fi
\fi
}%
}